# Unexpected Effects of Disorder on Current Fluctuations in the Symmetric Simple Exclusion Process

### v1

@@ VI Conclusion
In this paper, we have studied how the disorder affects the current fluctuations in the SSEP within a heterogeneous environment. First, we have provided the exact result (33) for the second moment of the integrated current for the SSEP with a defect site. The asymmetry in the dependence of the second moment of the integrated current on the density around $1/2$ arises from the differing dynamics of holes and particles. Surprisingly, in the low-density range, the second moment of the integrated current exceeds that of the homogeneous system. This discovery helps regulate both current fluctuations and particle diffusivity. Building on the insights gained from the SSEP with a defect site, we have derived the approximate expression (43) for the second moment of the integrated current for the SSEP on a quenched random energy landscape. The behavior of the second moment of the integrated current is greatly influenced by the energy landscape. Consequently, the presence of many-body effects induces sample-to-sample fluctuations in the fluctuation of the integrated current, even when single-particle dynamics do not exhibit such fluctuations.

### v2

@@ VI Conclusion
In this paper, we have studied how the disorder affects the current fluctuations in the SSEP within a heterogeneous environment. First, we have provided the exact result (33) for the second moment of the integrated current for the SSEP with a defect site. The asymmetry in the dependence of the second moment of the integrated current on the density around $1/2$ arises from the differing dynamics of holes and particles. Surprisingly, in the low-density range, the second moment of the integrated current exceeds that of the homogeneous system. Building on the insights gained from the SSEP with a defect site, we have derived the approximate expression (43) for the second moment of the integrated current for the SSEP on a quenched random energy landscape. The behavior of the second moment of the integrated current is greatly influenced by the energy landscape. Consequently, the presence of many-body effects induces sample-to-sample fluctuations in the fluctuation of the integrated current, even when single-particle dynamics do not exhibit such fluctuations. Our findings help control both current fluctuations and particle diffusivity by regulating the particle density. This approach to controlling particle behavior has potential applications in drug delivery systems and biological transport phenomena.